# GaussianFusion: Gaussian-Based Multi-Sensor Fusion for End-to-End Autonomous Driving

### v1

@@ Abstract
Multi-sensor fusion is crucial for improving the performance and robustness of end-to-end autonomous driving systems. Existing methods predominantly adopt either attention-based flatten fusion or bird’s eye view fusion through geometric transformations. However, these approaches often suffer from limited interpretability or dense computational overhead. In this paper, we introduce GaussianFusion, a Gaussian-based multi-sensor fusion framework for end-to-end autonomous driving. Our method employs intuitive and compact Gaussian representations as intermediate carriers to aggregate information from diverse sensors. Specifically, we initialize a set of 2D Gaussians uniformly across the driving scene, where each Gaussian is parameterized by physical attributes and equipped with explicit and implicit features. These Gaussians are progressively refined by integrating multi-modal features. The explicit features capture rich semantic and spatial information about the traffic scene, while the implicit features provide complementary cues beneficial for trajectory planning. To fully exploit rich spatial and semantic information in Gaussians, we design a cascade planning head that iteratively refines trajectory predictions through interactions with Gaussians. Extensive experiments on the NAVSIM and Bench2Drive benchmarks demonstrate the effectiveness and robustness of the proposed GaussianFusion framework. The source code will be released at https://github.com/Say2L/GaussianFusion.
@@ 1 Introduction
We evaluate GaussianFusion on the planning-oriented NAVSIM dataset [11]. Utilizing the ResNet-34 backbone—consistent with prior methods [7, 30, 13]—our approach achieves 85.0 EPDMS [26] and 88.9 PDMS [11], significantly surpassing current state-of-the-art methods. To further evaluate the generalization and robustness of our framework, we conduct experiments on the closed-loop benchmark Bench2Drive [22], where the results consistently demonstrate the effectiveness of GaussianFusion. The key contributions of this work are summarized as follows:
@@ 2 Related Work
BEV fusion methods [29, 21, 33, 28, 43] typically fuse multi-modal features within the BEV plane based on geometric projection relationships. ContFuse [29] employs 3D point clouds as intermediaries to project pixel-level image features onto the BEV plane for fusion. ThinkTwice [21] and DriveAdapter adopt the LSS framework [33] to transform image features into BEV space, which are then concatenated with LiDAR-derived BEV features. BEVFormer [28] and FusionAD [43] further extend the BEV representation by incorporating a height dimension into the BEV queries, enabling spatially-aware feature extraction from image inputs. This type of method addresses the interpretability and 3D modeling limitations inherent in flatten fusion methods. However, since operating on a dense BEV grid, these methods often incur significant computational and memory overhead.
In contrast to existing fusion paradigms, we propose GaussianFusion, which leverages Gaussian representations as intermediaries to extract and aggregate multi-modal features. Gaussians are well-regarded for their powerful scene reconstruction capabilities [25], and their physical attributes offer intuitive interpretability for researchers. Moreover, Gaussian representations are sparsely distributed in space, in stark contrast to the dense grids used in BEV fusion methods. As a result, our GaussianFusion effectively circumvents the limitations of both flatten fusion (lack of structure and interpretability) and BEV fusion (computational and memory inefficiency), offering a more efficient and interpretable alternative for multi-sensor integration.
@@ 3 Method
The objective of E2E autonomous driving is to predict the future trajectory of the ego vehicle directly from raw sensor inputs. Formally, given multi-view images =$\mathcal{I}=\{\mathbf{I}_{i}\in\mathbb{R}^{3\times H\times W}\mid i=1,\ldots,N\}$, LiDAR points $P\in\mathbb{R}^{n\times 3}$ and transformation matrices $\mathcal{T}=\{\mathbf{T}_{i}\in\mathbb{R}^{3\times 3}\mid i=1,\ldots,N\}$, the goal is to predict the ego vehicle trajectory $\tau=\{\left(x_{t},y_{t}\right)\}_{t=1}^{T}$, where $N$, $(H,W)$, $n$, $(x_{t},y_{t})$ and $T$ denotes the number of views, the image resolution, the number of points, the waypoint coordinate at time $t$ and the planning horizon, respectively.
@@ 3.3 Gaussian Decoder: Gaussians to Scene Reasoning
where $\operatorname{CrossAttn}$ denotes a cross-attention layer, and $\operatorname{Embedding}$ represents an embedding layer that encodes the anchor trajectory $\mathbf{A}$ into an initial query feature $\mathbf{F}_{\text{query}}$. The terms $\mathbf{f}^{\text{exp}}_{i}$ and $\mathbf{f}^{\text{imp}}_{i}$ refer to the explicit and implicit features of the Gaussian $\mathbf{G}_{i}\in\mathcal{G}_{A}$, respectively, and $[\mathbf{f}^{\text{exp}}_{i};\mathbf{f}^{\text{imp}}_{i}]$ denotes their concatenation. The traffic map and surrounding agents can be decoded from the most recent Gaussians $\mathcal{G}$, enabling $\mathcal{G}$ to serve as a comprehensive representation of the traffic scene. Therefore, we employ a cross-attention layer to build interactions between the anchor feature $\mathbf{F}_{A}$ and hidden features of $\mathcal{G}$. The refined trajectory $\tau=\{\left(x_{t},y_{t}\right)\}_{t=1}^{T}$ is obtained as follows:
@@
Table 1: Performance on the NAVSIM navtest split with extended metrics. ‘*’ represents that the results are sourced from Hydra-MDP++ [26]. Definitions of sub-metrics including ‘NC’, ‘DAC’, ‘EP’, ‘TTC’, ‘DDC’, ‘LK’, and ‘EC’ are provided in Appendix B. The best results are highlighted in bold.
@@ 5 Implementation Details
For the NAVSIM benchmark, we use the NAVSIM train split for training. For the Bench2Drive benchmark, the training data follows the TF++ [47] under the “12 only” setting. We utilize input from front, left-front, and right-front cameras, along with LiDAR point clouds. The camera images are cropped to a resolution of $448\times 250$. LiDAR points are projected onto the BEV plane, following the approach in TransFuser [7]. For a fair comparison with previous works [11, 30], we employ ResNet-34 [14] as the backbone. In our main experiments, the number of Gaussians is set to 512, and each Gaussian feature has a dimensionality of 128. We adopt 4 GaussianEncoder blocks and 2 cascade planning blocks. The number of anchor trajectories is set to 20 following [30]. It is worth noting that the map construction module in the GaussianDecoder is detached during inference for efficiency. Training is performed using the AdamW optimizer [32], with 50 epochs, a weight decay of $1\times 10^{-4}$, and a maximum learning rate of $6\times 10^{-4}$, which follows a cosine annealing schedule for learning rate decay. Hyper-parameter analysis is in Appendix C.
@@ 5.1 Comparison with State-of-the-Art Methods
Results on NAVSIM. We benchmark GaussianFusion against leading state-of-the-art (SOTA) methods on the NAVSIM navtest split. To ensure fair comparison, all models use ResNet-34 [14] as the backbone. As shown in Table 1, our approach achieves 85.0 EPDMS, outperforming recent strong baselines such as [30] and [13] by 1.0 and 1.9 points, respectively. Note that EPDMS poses a stricter challenge than PDMS by incorporating more nuanced driving criteria. A closer look reveals that most gains come from the Drivable Area Compliance (DAC) and Lane Keeping (LK) sub-metrics, indicating that GaussianFusion enables more stable and context-aware behavior in complex environments. Additionally, We evaluate performance under the standard PDMS metric, with results provided in Table 4 in the Appendix. Again, our method significantly surpasses previous methods, particularly on critical sub-metrics like DAC and Ego Vehicle Progress (EP). These results consistently affirm the robustness and effectiveness of our method across multiple evaluation protocols.
Results on Bench2Drive. We further conduct experiments on the closed-loop benchmark, Bench2Drive, to compare our method with existing SOTA E2E methods. As shown in Table 2, our method, GaussianFusion, achieves the best overall performance (79.4 DS), outperforming all learning-based baselines. It shows balanced strength across diverse tasks, with notable gains in overtaking and traffic sign compliance. Compared to the rule-based privileged method PDM-Lite, our method still falls short to some extent, indicating that there remains substantial room for improvement in E2E autonomous driving methods.
@@ 5.2 Ablation Study
Effect of Different Components. To understand the impact of each design choice in GaussianFusion, we perform a controlled ablation study by incrementally adding Gaussian Explicit Fusion (Gaussian Exp. Fusion), Gaussian Implicit Fusion (Gaussian Imp. Fusion), the Cascade Planning head (Cascade Planning), and the Agent Prediction head (Agent Pred.) to the TransFuser baseline [7]. The results are shown in Table 3. Introducing Gaussian Exp. Fusion leads to a substantial +2.4 gain in EPDMS, while also reducing the parameter count. Adding Gaussian Imp. Fusion further improves performance to 84.5 EPDMS, with only a slight increase in parameters. Furthermore, the inclusion of the Cascade Planning head pushes the score to 85.0 EPDMS, setting a new record with a total parameter count comparable to that of the baseline. These findings confirm that our architectural components offer strong performance gains without significantly increasing model complexity. Finally, we observe that incorporating the agent prediction head degrades performance. We attribute this to the agent prediction task failing to provide effective guidance for the Gaussian refinement process, instead introducing ambiguity—what we refer to as Gaussian confusion. Given that the semantic map already encodes sufficient agent-related information, we remove the agent prediction head from our final design.
Different Multi-Sensor Fusion Methods. Table 3 presents a comprehensive comparison of various multi-sensor fusion methods in terms of model parameters, semantic map construction, trajectory planning, and inference latency. The latency is measured by an RTX3090. To ensure a fair comparison, all methods adopt the same backbone and task heads. Our proposed method, GaussianFusion, achieves the best performance in both semantic map construction and trajectory planning while utilizing the fewest parameters, highlighting the effectiveness of our fusion strategy and the promise of Gaussian representations for autonomous driving.
@@
Benefiting from the sparsity of Gaussian representations, GaussianFusion also outperforms dense BEV fusion methods in terms of inference speed. However, its latency remains higher than that of the flatten fusion method, despite using fewer parameters. We attribute this to the presence of customized CUDA operations in our implementation, which are currently less optimized than native PyTorch operations.
@@ Appendix A Map Construction
[…]
@@
where $p_{\mathrm{b}}(\mathbf{x})$ and $p_{\mathrm{f}}(\mathbf{x})$ denote the background and foreground probabilities of pixel $\mathbf{x}$, $\operatorname{Softmax}(\cdot)$ represents the Softmax function. Intuitively, for a background pixel $\mathbf{x}_{\mathrm{b}}$, the $p_{\mathrm{b}}(\mathbf{x}_{\mathrm{b}})$ will decrease with Gaussians gather on $\mathbf{x}_{\mathrm{b}}$ because $p_{\mathrm{b}}(\mathbf{x}_{\mathrm{b}})\leq(1-\alpha(\mathbf{x}_{\mathrm{b}};\mathbf{G}_{i}))$ holds for any Gaussian. Conversely, for a foreground pixel $\mathbf{x}_{f}$, the $p_{\mathrm{f}}(\mathbf{x}_{\mathrm{f}})$ will increase with Gaussians gather on $\mathbf{x}_{\mathrm{f}}$ due to $1-p_{\mathrm{b}}(\mathbf{x}_{\mathrm{f}})\geq\alpha(\mathbf{x}_{\mathrm{f}};\mathbf{G}_{i})$ holds for any Gaussian. Therefore, Eq. 10 will encourage Gaussians to move to foreground pixels, as demonstrated in Figure 4. Given background and foreground predictions $[p_{\mathrm{b}};p_{\mathrm{f}}]\in\mathbb{R}^{C+1}$ and ground-truth bev map, we employ cross entropy loss $L_{ce}$ and the lovasz-softmax [2] loss $L_{lov}$ as the map construction loss functions.
Method Input NC $\uparrow$ DAC $\uparrow$ TTC $\uparrow$ EP $\uparrow$ PDMS $\uparrow$
LTF [7] Camera 97.4 92.8 92.4 79.0 83.8
TransFuser [7] C & L 97.7 92.8 93.0 79.2 84.0
GoalFlow* [41] C & L 98.3 93.3 94.8 79.8 85.7
Hydra-MDP++ [26] C & L 98.3 96.0 94.6 78.7 86.5
ARTEMIS [13] C & L 98.3 95.1 94.3 81.4 87.0
DiffusionDrive [30] C & L 98.2 96.2 94.7 82.2 88.1
GaussianFusion (Ours) C & L 98.3 97.2 94.6 83.0 88.8
Table 4: Performance on the Navtest Benchmark with Original Metrics.

### v2

@@ Abstract
Multi-sensor fusion is crucial for improving the performance and robustness of end-to-end autonomous driving systems. Existing methods predominantly adopt either attention-based flatten fusion or bird’s eye view fusion through geometric transformations. However, these approaches often suffer from limited interpretability or dense computational overhead. In this paper, we introduce GaussianFusion, a Gaussian-based multi-sensor fusion framework for end-to-end autonomous driving. Our method employs explicit and compact Gaussian representations as intermediate carriers to aggregate information from diverse sensors. Specifically, we initialize a set of 2D Gaussians uniformly across the driving scene, where each Gaussian is parameterized by physical attributes and equipped with explicit and implicit features. These Gaussians are progressively refined by integrating multi-modal features. The explicit features capture rich semantic and spatial information about the traffic scene, while the implicit features provide complementary cues beneficial for trajectory planning. To fully exploit rich spatial and semantic information in Gaussians, we design a cascade planning head that iteratively refines trajectory predictions through interactions with Gaussians. Extensive experiments on the NAVSIM and Bench2Drive benchmarks demonstrate the effectiveness and robustness of the proposed GaussianFusion framework. The source code will be released at https://github.com/Say2L/GaussianFusion.
@@ 1 Introduction
We evaluate GaussianFusion on the planning-oriented NAVSIM dataset [11]. Utilizing the V2-99 backbone [27], our approach achieves 92.0 PDMS [11], significantly surpassing current state-of-the-art methods. To further evaluate the generalization and robustness of our framework, we conduct experiments on the closed-loop benchmark Bench2Drive [23], where the results consistently demonstrate the effectiveness of GaussianFusion. The key contributions of this work are summarized as follows:
@@ 2 Related Work
BEV fusion methods [31, 22, 35, 30, 45] typically fuse multi-modal features within the BEV plane based on geometric projection relationships. MILE [15] lifts image features into 3D using predicted depth distributions, then aggregates voxel features into BEV via grid-based sum-pooling. ContFuse [31] employs 3D point clouds as intermediaries to project pixel-level image features onto the BEV plane for fusion. ThinkTwice [22] and DriveAdapter [21] adopt the LSS framework [35] to transform image features into BEV space, which are then concatenated with LiDAR-derived BEV features. BEVFormer [30] and FusionAD [45] further extend the BEV representation by incorporating a height dimension into the BEV queries, enabling spatially-aware feature extraction from image inputs. This type of method addresses the interpretability and 3D modeling limitations inherent in flatten fusion methods. However, since operating on a dense BEV grid, these methods often incur significant computational and memory overhead.
In contrast to existing fusion paradigms, we propose GaussianFusion, which leverages Gaussian representations as intermediaries to extract and aggregate multi-modal features. Gaussians are well-regarded for their strong scene reconstruction [26] and explicit representation properties. Moreover, Gaussian representations are sparsely distributed in space, in stark contrast to the dense grids used in BEV fusion methods. As a result, our GaussianFusion effectively circumvents the limitations of both flatten fusion (lack of structure and interpretability) and BEV fusion (computational and memory inefficiency), offering a more efficient and interpretable alternative for multi-sensor integration.
@@ 3 Method
The objective of E2E autonomous driving is to predict the future trajectory of the ego vehicle directly from raw sensor inputs. Formally, given multi-view images $\mathcal{I}=\{\mathbf{I}_{i}\in\mathbb{R}^{3\times H\times W}\mid i=1,\ldots,N\}$, LiDAR points $P\in\mathbb{R}^{n\times 3}$ and transformation matrices $\mathcal{T}=\{\mathbf{T}_{i}\in\mathbb{R}^{3\times 3}\mid i=1,\ldots,N\}$, the goal is to predict the ego vehicle trajectory $\tau=\{\left(x_{t},y_{t}\right)\}_{t=1}^{T}$, where $N$, $(H,W)$, $n$, $(x_{t},y_{t})$ and $T$ denotes the number of views, the image resolution, the number of points, the waypoint coordinate at time $t$ and the planning horizon, respectively.
@@ 3.3 Gaussian Decoder: Gaussians to Scene Reasoning
where $\operatorname{CrossAttn}$ denotes a cross-attention layer, and $\operatorname{Embedding}$ represents an embedding layer that encodes the anchor trajectory $\mathbf{A}$ into an initial query feature $\mathbf{F}_{\text{query}}$. The terms $\mathbf{f}^{\text{exp}}_{i}$ and $\mathbf{f}^{\text{imp}}_{i}$ refer to the explicit and implicit features of the Gaussian $\mathbf{G}_{i}\in\mathcal{G}_{A}$, respectively, and $\mathbf{f}^{\text{exp}}_{i};\mathbf{f}^{\text{imp}}_{i}]$ denotes their concatenation. The traffic map and surrounding agents can be decoded from the most recent Gaussians $\mathcal{G}$, enabling $\mathcal{G}$ to serve as a comprehensive representation of the traffic scene. Therefore, we employ a cross-attention layer to build interactions between the anchor feature $\mathbf{F}_{A}$ and hidden features of $\mathcal{G}$. The refined trajectory $\tau=\{\left(x_{t},y_{t}\right)\}_{t=1}^{T}$ is obtained as follows:
@@
Method Input Img. Backbone NC $\uparrow$ DAC $\uparrow$ TTC $\uparrow$ EP $\uparrow$ PDMS $\uparrow$
LTF Camera ResNet34 97.4 92.8 92.4 79.0 83.8
TransFuser C & L ResNet34 97.7 92.8 93.0 79.2 84.0
GoalFlow C & L ResNet34 98.3 93.3 94.8 79.8 85.7
Hydra-MDP C & L ResNet34 98.3 96.0 94.6 78.7 86.5
Hydra-MDP++ C ResNet34 97.6 96.0 93.1 80.4 86.6
ARTEMIS C & L ResNet34 98.3 95.1 94.3 81.4 87.0
DiffusionDrive C & L ResNet34 98.2 96.2 94.7 82.2 88.1
GaussianFusion (Ours) C & L ResNet34 98.3 97.2 94.6 83.0 88.8
GoalFlow C & L V2-99 98.4 98.3 94.6 85.0 90.3
Hydra-MDP-C C V2-99 98.7 98.2 95.0 86.5 91.0
Hydra-MDP++ C V2-99 98.6 98.6 95.1 85.7 91.0
GaussianFusion (Ours) C & L V2-99 98.7 98.1 95.7 88.2 92.0
Table 1: Performance on the Navtest Benchmark with Original Metrics. Definitions of sub-metrics are provided in Appendix B. The best results of different backbones are highlighted in bold, separately.
@@
Table 2: Performance on the NAVSIM navtest split with extended metrics. ‘*’ represents that the results are sourced from Hydra-MDP++ [28]. The best results are highlighted in bold.
@@ 5 Implementation Details
For the NAVSIM benchmark, we use the NAVSIM train split for training. We utilize input from front, left-front, and right-front cameras, along with LiDAR point clouds. The camera images are cropped to a resolution of $448\times 250$. For the Bench2Drive [23] benchmark, we adopt only the front-view camera, which already provides a wide field of view. In this case, we use the resolution of $1024\times 384$, which is consistent with the setting used in TransFuser++ [49]. For the two benchmarks, LiDAR points are projected onto the BEV plane, following the approach in TransFuser [7]. In our main experiments, the number of Gaussians is set to 512, and each Gaussian feature has a dimensionality of 128. We adopt 4 GaussianEncoder blocks and 2 cascade planning blocks. The number of anchor trajectories is set to 20 following [32]. It is worth noting that the map construction module in the GaussianDecoder is detached during inference for efficiency. Training is performed using the AdamW optimizer [34], with 50 epochs, a weight decay of $1\times 10^{-4}$, and a maximum learning rate of $6\times 10^{-4}$, which follows a cosine annealing schedule for learning rate decay. Hyper-parameter analysis is in Appendix C.
@@ 5.1 Comparison with State-of-the-Art Methods
Results on NAVSIM. We benchmark GaussianFusion against leading state-of-the-art (SOTA) methods on the NAVSIM navtest split. We employ ResNet34 [14] and V2-99 [27] as the image backbone of GaussianFusion, respectively. As shown in Table 1, our method significantly surpasses previous methods, particularly on critical sub-metrics like DAC and Ego Vehicle Progress (EP) under the ResNet34 setting. Our enhanced GaussianFusion, equipped with V2-99, further improves performance across all metrics—achieving state-of-the-art results on the Navtest benchmark. This demonstrates the scalability of our framework with stronger backbones. Additionally, we evaluate performance under the EPDMS metric. Note that EPDMS poses a stricter challenge than PDMS by incorporating more nuanced driving criteria. As shown in Table 2, our approach achieves 85.0 EPDMS, outperforming recent strong baselines such as [32] and [13] by 1.0 and 1.9 points, respectively. A closer look reveals that most gains come from the Drivable Area Compliance (DAC) and Lane Keeping (LK) sub-metrics, indicating that GaussianFusion enables more stable and context-aware behavior in complex environments. These results consistently affirm the robustness and effectiveness of our method across multiple evaluation protocols.
Results on Bench2Drive. We further conduct experiments on the closed-loop benchmark, Bench2Drive, to compare our method with existing SOTA E2E methods. Due to high performance variability in the CARLA simulator, we report results across three different random seeds. As shown in Table 3, our method, GaussianFusion, achieves the best overall performance (79.4 DS), outperforming all learning-based baselines. It shows balanced strength across diverse tasks, with notable gains in overtaking and traffic sign compliance. Compared to the rule-based privileged method PDM-Lite, our method still falls short to some extent, indicating that there remains substantial room for improvement in E2E autonomous driving methods.
@@ 5.2 Ablation Study
Effect of Different Components. To understand the impact of each design choice in GaussianFusion, we perform a controlled ablation study by incrementally adding Gaussian Explicit Fusion (Gaussian Exp. Fusion), Gaussian Implicit Fusion (Gaussian Imp. Fusion), the Cascade Planning head (Cascade Planning), and the Agent Prediction head (Agent Pred.) to the TransFuser baseline [7]. The results are shown in Table 4. Introducing Gaussian Exp. Fusion leads to a substantial +2.4 gain in EPDMS, while also reducing the parameter count. Adding Gaussian Imp. Fusion further improves performance to 84.5 EPDMS, with only a slight increase in parameters. Separately, incorporating Cascade Planning with only the explicit fusion also improves trajectory prediction, increasing LK and EPDMS compared to using explicit fusion alone. The best overall performance is achieved when combining dual-branch fusion and Cascade Planning, reaching an EPDMS of 85.0. These findings confirm that our architectural components offer strong performance gains without significantly increasing model complexity. Finally, we observe that incorporating the agent prediction head degrades performance. We attribute this to the agent prediction task failing to provide effective guidance for the Gaussian refinement process, instead introducing ambiguity—what we refer to as Gaussian confusion. Given that the semantic map already encodes sufficient agent-related information, we remove the agent prediction head from our final design.
Different Multi-Sensor Fusion Methods. Table 3 presents a comprehensive comparison of various multi-sensor fusion methods in terms of model parameters, semantic map construction, trajectory planning, and inference latency. The latency is measured by an RTX3090. To ensure a fair comparison, all methods adopt the same backbone and task heads.
Compared with the BEV fusion method, our Gaussian fusion achieves better performance (+0.04 mIoU and +0.7 EPDMS), while also significantly reducing computational cost: the FLOPs are 38.9% lower and latency is reduced by 25.6%, showing clear advantages in both accuracy and efficiency.
@@
Compared with the flatten (TransFuser-style) fusion, which is a very lightweight approach, our method exhibits only a 30.3% increase in FLOPs, while achieving much stronger performance (+0.05 mIoU and +2.7 EPDMS). However, our method’s latency is 77.8% higher, which we believe is largely due to the differences in operator-level optimization. This could be further mitigated through engineering improvements in deployment.
@@ 6 Conclusion
Limitations. We have not conducted quantitative evaluations on tracking small or fast-moving objects. Similarly, our method has not been assessed under sensor occlusion or noise, which are critical factors for real-world deployment. Moreover, our method depends on 3D bounding box and road topology labels to optimize the 2D Gaussians. While effective, such supervision can be expensive or unavailable in real-world scenarios.
Acknowledgments
This work was supported in part by the the Guangxi Key R & D Program under Grant No.GuikeAB24010324, and in part by the Guangdong Basic and Applied Basic Research Foundation under Grant 2025A1515011485.
@@ Appendix A Map Construction
where $\mathbf{m}_{i}$, $\mathbf{s}_{i}$, and $\mathbf{r}_{i}$ denote the mean, scale, and rotation vectors of the Gaussian $\mathbf{G}_{i}$, respectively. $\boldsymbol{\Sigma}_{i}$ represents the covariance matrix of $\mathbf{G}_{i}$; $\operatorname{diag}(\cdot)$ denotes the operator that constructs a diagonal matrix from a given vector; and $\operatorname{r2m}(\cdot)$ denotes the function that converts a rotation vector into its corresponding rotation matrix. Then the superposition probability can be calculated as:
@@
where $p_{\mathrm{b}}(\mathbf{x})$ and $p_{\mathrm{f}}(\mathbf{x})$ denote the background and foreground probabilities of pixel $\mathbf{x}$, $\operatorname{Softmax}(\cdot)$ represents the Softmax function. Intuitively, for a background pixel $\mathbf{x}_{\mathrm{b}}$, the $p_{\mathrm{b}}(\mathbf{x}_{\mathrm{b}})$ will decrease with Gaussians gather on $\mathbf{x}_{\mathrm{b}}$ because $p_{\mathrm{b}}(\mathbf{x}_{\mathrm{b}})\leq(1-\alpha(\mathbf{x}_{\mathrm{b}};\mathbf{G}_{i}))$ holds for any Gaussian. Conversely, for a foreground pixel $\mathbf{x}_{f}$, the $p_{\mathrm{f}}(\mathbf{x}_{\mathrm{f}})$ will increase with Gaussians gather on $\mathbf{x}_{\mathrm{f}}$ due to $1-p_{\mathrm{b}}(\mathbf{x}_{\mathrm{f}})\geq\alpha(\mathbf{x}_{\mathrm{f}};\mathbf{G}_{i})$ holds for any Gaussian. Therefore, Eq. 10 will encourage Gaussians to move to foreground pixels, as demonstrated in Figure 4. Given background and foreground predictions $p_{\mathrm{b}};p_{\mathrm{f}}]\in\mathbb{R}^{C+1}$ and ground-truth bev map, we employ cross entropy loss $L_{ce}$ and the lovasz-softmax [2] loss $L_{lov}$ as the map construction loss functions.